10 December 2013: Tuesday

LDAP:

  • No issues in last night’s update.
  • VPN flipping…
  • OSP code change testing, and production roll-out.
    • EDIT: 12/11 – Because the OSPEmail attribute does not have an EQUALITY match rule, the update code needs to be updated to replace instead of delete/add.  Took two (2) hours (8:30-10:30) last night to manually process the rest of the update.

Backups:

  • Cleaning cartridge expired at 0525.  Replaced at 0649.
  • No response from EMC about the open issue (59521968).  Need to poke them again.  Scheduled a call for 10am tomorrow.
  • Windows disaster recovery group had three failures.  This is getting to be a pattern.  Now, if I could only figure out what the pattern indicates :/
  • df -h reports 2.3TB left on that second ZFS pool that I’m trying to empty off.  Oops!  I forgot about the gzip-1 compression.  It’s really over 7TB of data.  Well, darn!  That’s going to take a while to stage off.  Perhaps my Christmas present to myself will be getting this done before New Years??? (or maybe not)
  • recover from over the summer required.  Two tapes retrieved from the vaults.  Recover command instructions given to the IT support person.

FootPrints:

  • Nothing back from Numara/BMC – perhaps it is time I gave them a phone call?

General:

  • Lots of work trying to adjust to the new firewall rule check in cfengine.

Tags:

Leave a Reply

You must be logged in to post a comment.